OpenStudy (anonymous):

How many different arrangements of the letters in the word BETTER are there?

OpenStudy (anonymous):

7!/3!

OpenStudy (anonymous):

4^6=4096

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

haven't seen the right answer yet

OpenStudy (anonymous):

isn't it 4096?

OpenStudy (anonymous):

the choices are 180,360, and 720

OpenStudy (anonymous):

if the letters were all different, it would be \(6!\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

but since you cannot tell apart the two "E" and two "T" it is \[\frac{6!}{2\times 2}\]

OpenStudy (anonymous):

but the letters have multiples

OpenStudy (anonymous):

360

OpenStudy (anonymous):

4096?

OpenStudy (anonymous):

\[\frac{6\times 5\times 4\times 3\times 2}{4}=6\times 5\times 3\times 2\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i bet

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if all the letters were different (they are not, but i am pretending) there would be 6 choices for the first letter 5 for the second 4 for the third 3 for the fourth 2 for the fifth and 1 for the sixth

OpenStudy (anonymous):

by the counting principle the number of ways to do this is found by multiplying all the choices together, i.e. \[6\times 5\times 4\times 3\times 2\times 1=6!\]

OpenStudy (anonymous):

so it is 360

OpenStudy (anonymous):

but in this case you have repeated letters there are two E and two T so you have overcounted

OpenStudy (anonymous):

you have to divide by 2 twice, because of this

OpenStudy (anonymous):

why do u divide by 2 twice?

OpenStudy (anonymous):

so the answer is \[\frac{6\times 5\times 4\times 3\times 2}{4}=6\times 5\times 3\times 2\]

OpenStudy (anonymous):

because there are two E and also two T

OpenStudy (anonymous):

i get 180

OpenStudy (anonymous):

i get all of it except why divide by 4

OpenStudy (anonymous):

could u go into depth

OpenStudy (anonymous):

because for example \[E_1BE_2T_1T_2R\] looks the same as \[E_2BE_1T_1T_2R\] looks the same as \[E_1BE_2T_2T_1R\] looks the same as \[E_2BE_1T_2T_1R\] without the subscripts

OpenStudy (anonymous):

the \(6!\) counts these as different, but in fact they are the same \(6!\) overcounts by a factor of 4

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

yw